OpenStudy (anonymous):

The pyramids at Giza include several square-based pyramids that were built in honor of the Egyptian gods. One of these incredible monuments is the Khafre pyramid, built as part of a temple complex, with the famous Sphinx at its base. The base length of this pyramid is seven hundred five feet, and its height is four hundred seventy-one feet. Massive dimensions indeed! Calculate the volume of the Khafre pyramid. Show the steps of your solution and explain your work.

OpenStudy (anonymous):

@mtbender74 can you help me with this question?

OpenStudy (anonymous):

let me take a look at it

OpenStudy (anonymous):

ok...so the volume of a pyramid is V = 1/3 * (area of the base) * h (looks familiar?) the base is a square with side length 705 and the height is 471.

OpenStudy (anonymous):

can you take it from there?

OpenStudy (anonymous):

hold on let me think about this

OpenStudy (anonymous):

so the base is 705 so its v = 1/3 * 705 * 471 correct?

OpenStudy (anonymous):

almost...the base length is 705...but the area of a square is s^2. so you need to square the 705

OpenStudy (anonymous):

how do i do0 that

OpenStudy (anonymous):

705 * 705

OpenStudy (anonymous):

so 497025

OpenStudy (anonymous):

that's the area of the base...now put it into the formula

OpenStudy (anonymous):

the number is going to be big :)

OpenStudy (anonymous):

so now it would be V = 1/3 * 497025 * 471

OpenStudy (anonymous):

which equals?

OpenStudy (anonymous):

well then it would go V = 1/3 * 2.3409877.... i think this is wrong..

OpenStudy (anonymous):

497025 * 471 = 234098775... divide this by 3 (which is the same as multiplying by 1/3)...

OpenStudy (anonymous):

78032925?

OpenStudy (anonymous):

looks good

OpenStudy (anonymous):

okay so thats the final answer? 78032925

OpenStudy (anonymous):

cubic ft., yes